OpenStudy (anonymous):

6(p)^2 - 2(p) + 3 = 3

OpenStudy (anonymous):

Subtract 3 from both the sides, let us see what you get..

OpenStudy (anonymous):

= 0. Already have my answer. Need to double check with someone. (: Am confused. ):

OpenStudy (anonymous):

What is your answer??

OpenStudy (anonymous):

\[6p^2−2p=3−3=0\] \[2p(3p−1)=0\] 2p = 0 and (3p - 1) = 0 p = 0 and p = 1/3 is the answer...